A crane is lifting a 123 kg beam with an acceleration of 1.26 m/s2 downwards. Calculate the magnitude of the tension in the cable, to 2 sf.
Answer #1

Given

crane lifting mass is m = 123 kg , with an acceleration of a = 1.26 m/s^2 downwards

Tension T acts upwards and the force of gravity acts downward and the acceleration also downward

it can be written as T - mg = -ma

T = m(g-a)

T = 123(9.8-1.26) N

T = 1050.42 N

so the magnitude of the tension in the cable is T = 1050.42 N